Solve the equation for the missing variables
1. 12 (X-3)= 0
2. 15 (x+1)-8x= 29
3. -8 (x-6)= 3(x-6)
4. -3 (5x + 2)+ 5(1+3x)= 0
5. 7=3(x+2)-3(x-5)
6. 74-12x= 2

Answer by Edwin McCravy(20059)

We are only supposed to do two.  So I'll do 3 and 4:

3.    -8(x-6) = 3(x-6)

Didtribute on both sides:

       -8x+48 = 3x-18

Subtract 3x from both sides:

      -11x+48 = -18

Subtract 48 from both sides:

         -11x = -66

            x = 6

========================

4.   -3(5x+2) + 5(1+3x) = 0

Distribute:

           -15x-6+5+15x = 0

Combine like terms (the 15x and -15x cancel)

                     -1 = 0

This is a false numerical equation, so there
is no solution.

Edwin